# HELM: Hierarchical Encoding for mRNA Language Modeling

### v1

@@ 4.4 HELM improves Ab-mRNA sequence region annotation
Table 4 shows that HELM significantly improves annotation accuracy compared to the non-hierarchical XE model, particularly when using the Masked Language Modeling (MLM) objective. This improvement can be attributed to two factors. First, HELM’s hierarchical prior better captures the inherent structure of mRNA sequences, where certain regions are functionally and structurally dependent on others. Second, the bidirectional attention in MLM more effectively captures the context needed for precise sequence annotation, which is crucial for identifying distinct regions where both upstream and downstream nucleotide dependencies are important. For details see (Sec. A.8).
@@ A.4.2 Results
A.5 Entropy Details
@@
Table 7: Average entropy values of synonymous codon distributions for the top five most frequent amino acids in each dataset. Datasets with lower entropy values, such as MLOS, Tc-Riboswitches, mRFP, and iCodon, exhibit stronger synonymous codon usage bias, correlating with the enhanced performance of hierarchical HELM models on these datasets (see Fig. 3 and Table 2).
@@
Ab1 1.81
Ab2 1.81
MLOS 1.62
iCodon 1.77
Tc-Riboswitches 1.63
mRFP 1.20
A.6 Impact of the choice of $\alpha$ for HXE loss
@@
A.7 Scalability Experiments
In this section, we provide the details of our experiments with three GPT-2-based models: 20 million (20M), 50 million (50M), and 100 million (100M) parameters. All models were trained using the HELM framework with a Causal Language Modeling (CLM) objective. The performance was evaluated on downstream mRNA property prediction tasks.
A.7.1 Model Configurations
  • 20M Model: This model consists of 12 transformer layers, each with hidden size of 384 and intermediate size of 1536.
@@ A.7.1 Model Configurations
All models were trained on the same mRNA dataset using the HELM-modified CLM objective, and we maintained a consistent training pipeline across all experiments.
A.7.2 Results
Figure 7: Performance comparison of GPT-2 models with 20M, 50M, and 100M parameters trained using the HELM framework.
This analysis illustrates the scalability of HELM and highlights the balance between model size and performance efficiency (see Fig.  7. The diminishing returns beyond 50M parameters suggest that optimal model sizes for mRNA tasks may lie within the 50M-100M parameter range.
The training duration for each model configuration is illustrated in Figure 8.
Figure 8: Average Pre-training Time Across Models with Varying Parameter Counts
A.8 Ab-mRNA sequence region annotation experiment
@@
A.9 Pre-training data curation and statistics
Figure 9: Comparison of gene distributions between the original OAS human data and the curated dataset. (a) shows the gene distribution for heavy chains, and (b) shows the gene distribution for light chains. (c) illustrates the balanced numbers of heavy and light chains, including the distribution of Kappa and Lambda isotypes, along with V-allele distribution.

### v2

@@ 4.4 HELM improves Ab-mRNA sequence region annotation
Appendix Table 9 shows that HELM significantly improves annotation accuracy compared to the non-hierarchical XE model, particularly when using the Masked Language Modeling (MLM) objective. This improvement can be attributed to two factors. First, HELM’s hierarchical prior better captures the inherent structure of mRNA sequences, where certain regions are functionally and structurally dependent on others. Second, the bidirectional attention in MLM more effectively captures the context needed for precise sequence annotation, which is crucial for identifying distinct regions where both upstream and downstream nucleotide dependencies are important. For details see (Sec. A.9).
@@ 5 Conclusions
Reproducibility Statement
We are committed to the transparency and reproducibility of our findings. Code, datasets, and model weights will be made publicly available to the research community, allowing others to reproduce, verify, and extend our work. Additionally, to ensure reproducibility, we have also provide detailed descriptions of model architecture, hyperparameters, training and inference procedures in Appendix Sections A.2 and A.3 with additional details on experimental information.
@@ A.4.2 Results
A.5 Codon Usage Bias Analysis
A.5.1 Entropic codon bias
@@
Table 5: Average entropy values of synonymous codon distributions for the top five most frequent amino acids in each dataset reveal varying degrees of codon usage bias. Datasets like MLOS, Tc-Riboswitches, and mRFP, with lower entropy values, exhibit stronger codon usage bias, correlating with the enhanced performance of hierarchical HELM models. In contrast, datasets like iCodon and COV-19 Vaccine show relatively weaker codon usage bias, with Ab1 and Ab2 exhibiting the least bias, aligning with the comparatively smaller yet significant performance gains of hierarchical HELM models on these datasets (see Fig. 3 and Table 2).
@@
Ab1 1.80
Ab2 1.93
MLOS 1.28
iCodon 1.93
COV-19 Vaccine 2.07
Tc-Riboswitches 1.73
mRFP 1.37
A.6 Evaluating hierarchical encoding through synonymous sequence clustering
MLM Silh. ($\uparrow$) CLM Silh. ($\uparrow$)
XE 0.74 0.68
HELM 0.91 0.83
A.7 Impact of the choice of $\alpha$ for HXE loss
@@
A.8 Is learning hierarchy beneficial for larger scale models?
In this section, we provide the details of our experiments with two scales of transformer-based models: 50 million (50M), and 100 million (100M) parameters. All models were trained using the non-hierarchical XE and our hierarchical HELM framework with a Causal Language Modeling (CLM) objective. The performance was evaluated on downstream mRNA property prediction tasks on five datasets introduced in main text.
A.8.1 Model Configurations
@@
All models were trained on the same mRNA dataset using the either XE or HELM-modified CLM objective, and we maintain a consistent training pipeline across all experiments.
A.8.2 Encoding hierarchy helps even with scaled up models
As shown in Appendix Table 8, the 100M HELM model still significantly outperforms its XE counterpart, maintaining the relative performance gap observed at the 50M scale. This confirms that XE models do not naturally capture hierarchical information, even with increased size, unless explicitly guided by hierarchical cross-entropy (HXE) loss. While further scaling to even larger models would provide additional insights, our computational constraints (pretraining time of  2 weeks for 100M models) highlight the practical advantage of HELM’s hierarchical priors in enabling superior performance without requiring expensive scaling.
Table 8: Performance comparison of HELM vs. XE models for 50M and 100M parameter sizes. Spearman rank correlation is reported with bold indicating the best performance in each case.
Model Size Ab1 Ab2 MLOS
50M XE 0.752 0.597 0.611
50M HELM 0.760 0.614 0.592
100M XE 0.750 0.596 0.608
100M HELM 0.765 0.619 0.664
Note that even with larger scales, encoding hierarchy is useful. However, scaling the model incurs longer pre-training times with 50M and 100M models taking approximately 49.5 hours and 82 hours respectively.
A.9 Ab-mRNA sequence region annotation experiment
@@
MLM Acc. ($\uparrow$) CLM Acc. ($\uparrow$)
XE 78.68 67.51
HELM 83.39 73.39
A.10 Property prediction results in terms of other metrics
While we have reported property prediction performance in the main text in terms of the commonly used Spearman rank correlation metric, here we quantify results also in terms of Pearson correlation and $R^{2}$, see Table 10. Similar to the trend observed with Spearman correlation, our proposed HELM model outperforms baselines for all datasets in terms of these metrics as well.
Table 10: Performance comparison of HELM vs. XE models using Pearson and $R^{2}$ metrics. HELM models encoding mRNA hierarchy outperform non-hierarchical XE models across all downstream tasks and datasets, highlighting the importance of hierarchy as a strong biological prior. Bold indicates the best performing model.
Model Ab1 Ab2 MLOS Tc-Riboswitches mRFP COV-19 Vaccine
Transformer XE (MLM) 0.766 / 0.587 0.599 / 0.361 0.667 / 0.444 0.532 / 0.283 0.779 / 0.606 0.788 / 0.620
Transformer HELM (MLM) 0.793 / 0.629 0.603 / 0.364 0.719 / 0.516 0.601 / 0.361 0.847 / 0.717 0.811 / 0.657
Transformer XE (CLM) 0.757 / 0.573 0.596/ 0.356 0.563 / 0.316 0.485 / 0.235 0.880 / 0.774 0.762 / 0.580
Transformer HELM (CLM) 0.787 / 0.619 0.608 / 0.370 0.542 / 0.293 0.556 / 0.309 0.886 / 0.784 0.769 / 0.591
A.11 Analysis of HELM vs XE with different sequence lengths and GC content
Sequence Length Analysis
In this experiment, we divide the iCodon thermostability dataset into three sequence length categories: 30-1000, 1000-2000, and 2000-3000 nucleotides. These categories help analyze how model performance varies as sequence length increases. For mRNA vaccines, the typical length of sequences falls within the range of 1000-1500 nucleotides (Gunter et al., 2023). Thus, our analysis of sequences in the 2000-3000 nucleotide range covers longer sequences than typically observed in mRNA vaccine datasets.
For both non-hierarchical XE and hierarchical HELM models, performance decreases as the sequence length increases as shown in Appendix Table 11. This drop is due to the fact that our pre-training dataset includes sequences with a maximum length of approximately 1400 nucleotides. Hence, sequences longer than this range are less effectively represented by both XE and HELM models.
Despite the performance degradation, HELM models exhibit much less performance drop compared to XE models, particularly in the longer sequence categories as exhibited by Appendix Table 11. This suggests that HELM’s hierarchical encoding mechanism helps maintain performance even with longer sequences.
Table 11: Thermostability prediction performance of XE and HELM models across different sequence length ranges shown with Spearman rank correlation (higher value indicates better performance). The performance of both models decreases with increasing sequence length, particularly in the 2000-3000 nucleotide range. HELM demonstrates more robustness to sequence length variation.
Sequence Length Range XE (CLM) HELM (CLM) Number of Sequences
30-1000 0.523 0.530 23929
1000-2000 0.496 0.528 28943
2000-3000 0.468 0.514 12484
Sequence Length Range XE (MLM) HELM (MLM) Number of Sequences
30-1000 0.520 0.531 23929
1000-2000 0.506 0.529 28943
2000-3000 0.475 0.506 12484
GC Content Analysis
We also categorize the iCodon thermostability dataset based on GC content into three ranges: GC $\leq$ 47, 47 $<$ GC $\leq$ 55, and GC $>$ 55. This classification closely aligns with widely recognized thresholds in the literature, where GC content below 45% is considered low, and above 56% is high.
Performance for both XE and HELM decreases with extremely high GC contents as shown in Appendix Table 12. Again, this is likely due to the fact that pre-training data mostly contains sequences with normal GC content range (GC $<$ 50) and higher GC content sequences are less common. Once again, HELM models show superior performance compared to XE models across all GC content ranges, particularly for the high GC content range, where XE shows more pronounced performance drop than HELM.
Table 12: Thermostability prediction performance of XE and HELM models across different GC content ranges shown with Spearman rank correlation (higher value indicates better performance). Performance decreases for both models with higher GC content, but HELM remains more robust.
GC Content Range XE (CLM) HELM (CLM) Number of Sequences
GC $\leq$ 47 0.568 0.562 21783
47 $<$ GC $\leq$ 55 0.440 0.485 21803
GC $>$ 55 0.517 0.529 21770
GC Content Range XE (MLM) HELM (MLM) Number of Sequences
GC $\leq$ 47 0.557 0.570 21783
47 $<$ GC $\leq$ 55 0.446 0.496 21803
GC $>$ 55 0.516 0.524 21770
A.12 Additional baseline
Table 13: Performance comparison of HELM models with hierarchical cross-entropy (HXE) loss vs. baseline models with separate codon-type embeddings using XE loss. Spearman rank correlation is reported. Bold indicates the best performance.
Model Ab1 Ab2 MLOS Tc-Riboswitches mRFP COV-19 Vaccine
Baseline Sep. codon embedding (MLM) 0.736 0.579 0.457 0.572 0.810 0.780
HELM (MLM) 0.767 0.609 0.701 0.626 0.822 0.833
Baseline Sep. codon embedding (CLM) 0.729 0.572 0.521 0.599 0.825 0.781
HELM (CLM) 0.760 0.614 0.592 0.619 0.849 0.789
We also implement a simpler baseline by introducing separate embeddings for different types of codons (e.g., [start], [stop], [S], [Q]) and combining these with the token embeddings as input, while using the standard cross-entropy (XE) loss without explicitly modeling the hierarchy. We train both MLM and CLM models with this setup and evaluated their performance on downstream property prediction datasets introduced in main text. The results, reported in Table 13, demonstrate that HELM models, which explicitly incorporate hierarchy, consistently outperform this baseline on all datasets irrespective of the pre-training strategy (MLM or CLM). This further highlights the importance of hierarchical modeling in achieving robust performance across diverse datasets.
A.13 Generalization of models with Ab pre-training to general mRNA sequences
As reported in the main text (Table 2), both XE and HELM models generalize well to various types of mRNA sequences even though they have been pre-trained on the Ab-mRNA dataset. This generalization can be attributed to two key mechanisms:
  • Conserved RNA structural features: RNA molecules, whether synthetic or natural, share conserved structural motifs like loops, bulges, and stems that are critical for stability and function (Saito & Inoue, 2009).
  • Shared molecular interactions: as antibody mRNAs often encode features related to interactions with nucleic acid-like antigens, creating overlap in sequence characteristics with other types of RNA molecules  (Alberts et al., 2002).
These factors contribute to generalization of features learned from Antibody-mRNA to other types of mRNA observed in our experiments.
A.14 Pre-training data curation and statistics